Test, May 4,2018: 9-12 nidental usp n id siur on eick and tisiluminaled by white lghth is oil slick (i.e, determine the most constructively reflected wavelength from th 459 to its surface. What color will the oil slick appear to an observer looking down at the 14,n 1.33. White light has wavelength range 400 nm -700 nm.
Answer #1

for oil film, the equation for constructive interference is given by:

2*n_oil*d*cos(theta)=(m-0.5)*lambda

where n_oil=refractive index of oil=1.4

d=thickness of oil slick=150 nm

theta=incident angle=45 degrees

m is any integer

hence lambda=2*n_oil*d*cos(theta)/(m-0.5)

for different values of m,

lambda values are given as :

m   lambda

1   594 nm

2    198 nm

3   118.8 nm

so for m=1, wavelength falls between the wavelength range of 400 nm to 700 nm

hence corresponding color is orange.